# Exocentric To Egocentric Transfer For Action Recognition: A Short Survey

### v1

@@ Abstract
Egocentric vision captures the scene from the point of view of the camera wearer while exocentric vision captures the overall scene context. Jointly modeling ego and exo views is crucial to developing next-generation AI agents. The community has regained interest in the field of egocentric vision. While the third-person view and first-person have been thoroughly investigated, very few works aim to study both synchronously. Exocentric videos contain many relevant signals that are transferrable to egocentric videos. In this paper, we provide a broad overview of works combining egocentric and exocentric visions.
Index Terms:
Exocentric, egocentric, ego-exo learning, action recognition.
I Introduction
Hbeings perceive the world from multiple viewpoints. We watch Do-it-yourself videos to learn new skills. A bicycle repair video alternates between the ego (first-person) and exo (third-person) viewpoints. An ego (close-up) view of the bicycle captures vital hand-object interactions and an exo (third-person) view captures the overall context in the environment. We can relate the object from the third-person to the first-person perspective. Being able to map skills to one’s own body has been a well-studied problem in cognitive science [1, 2, 3]. Capturing video from both the ego and exo views is a vital frontier for AI to understand human activities. Widespread applications exist in augmented reality [4, 5] and robotics [6, 7, 8].
Despite the importance of multi-view learning, most efforts into video understanding have focused to only one view, third-person (exocentric) viewpoint [9, 10, 11, 12, 13, 14] or first-person (egocentric) viewpoints [15] separately. While existing algorithms perform considerably well on third-person settings [16], a significant gap exists in the egocentric settings [17, 18, 19].
An exocentric view contains many relevant cues for recognizing the egocentric view. For example in Fig 1, the hand-object interaction of ”cutting” in the third-person view can be useful to recognize in the first-person view.
In recent years, Vision-Language Models (VLMs) have shown their effectiveness on various vision tasks, such as action anticipation [20], video anomaly detection [21], and affective understanding [22]. Existing VLMs [23, 24] trained on large amounts of third-person perspective contain many signals that can be transferable for egocentric tasks. Previous works like [25, 26] utilize simpler architectures to learn egocentric representations from third-person data. VLMs are more capable of learning stronger representations. This short survey provides a timely high-level overview of various egocentric-exocentric learning tasks.
@@ I Introduction
II Datasets
Several datasets containing paired ego-exo views have been proposed in the literature [29, 30, 31, 32, 33]. ”Mixed” ego-exo views have been covered by [34, 35, 36, 37, 38, 39, 40, 41, 42]. Zhang et al. [43] captures egocentric interactions in a 3D viewpoint. Xue et al.’s AE2 dataset [42] is sampled from existing ego and exo datasets. These datasets have several shortcomings: lack of magnitude, weak synchronization, and poor diversity. The release of two new large-scale datasets [17, 27, 28] attempts to bridge this gap. Refer to Table I for an in-depth review.
TABLE I: Comparison of existing datasets across various parameters, arranged in chronological order.
@@ II Datasets
The CMU-MMAC dataset [30] is one of the earliest datasets that captures ego and exo video. It is composed of 43 participants cooking 5 recipes in the kitchen setting. Multiple modalities like audio, video, accelerations, and motion capture are present in this dataset.
@@
The Charades-Ego dataset [33] was one of the former large-scale joint multi-view dataset efforts containing 68.8 hours of first and third-person video. 112 actors hired by Amazon Mechanical Turk recorded 34 hours of scripted scenarios.
@@
The Lemma dataset [44] is composed of multi-view and multi-agent daily-life activities. 3D skeletons and RGB-D are collected to give a broad perspective.
@@
The Homage dataset [31] is a synchronized multi-view dataset consisting of 30 hours of ego-exo video from 27 participants performing household activities in the same environment. It is well annotated with both the hierarchical and atomic action labels.
@@
The H2O dataset [29] focuses on 3D egocentric object-level manipulations. It is composed of 3D hand poses, 6D object poses, camera poses, object meshes, and scene-point clouds. 4 different participants perform 36 unique actions in three unique environments.
@@
The Assembly101 dataset [32] features non-scripted multi-step activities. 101 toy vehicles are manipulated in 4321 video sequences for a total of 513 hours. It constitutes 1380 fine-grained and 202 coarse-grained action classes. AssemblyHands [45] is a subset of Assembly101 to study the challenging problem of 3D hand pose estimation and action classification.
@@
The AE2 dataset [42] is one of the premier attempts to learn a view-invariant self-supervised embedding from unpaired ego and exo videos. To this end, they create a new benchmark, sampled from five public datasets [19, 30, 46, 29, 47], and a self-collected tennis dataset. It is composed of 322 clips.
@@
The Ego-Exo4D dataset [17] is the largest multi-view dataset including the egocentric view and the corresponding exocentric information. Moreover, it also offers multiple natural language descriptions including expert commentary, narrate-and-act descriptions, and atomic action descriptions. It is rich in modalities like audio, IMU, video, depth, gaze, stereo, 3D environments, thermal IR, GPS, motion capture, 6DOF, barometer and magnetometer readings. 740 subjects shot 123 scenes across different cities. It releases new challenging benchmarks like keystep recognition, efficient action detection, and proficiency estimation.
@@
The EgoExoLearn dataset [27] is another concurrent large-scale ego-exo synchronized dataset. It contains 120 hours of demonstration activities recorded in the lab and daily-life settings. It is richly annotated with fine-grained captions. Unlike previous datasets, it releases benchmarks on cross-view action anticipation and proficiency estimation.
@@
The EgoExo-Fitness dataset [28] was also concurrently released along with the previous two ego-exo datasets. While the previous datasets extensively explored daily-life activities, EgoExo-Fitness focuses on exercise-related activities. It comes with a new set of benchmarks for cross-view sequence verification.
III Existing Approaches
Figure 3: Overview of the landscape of various joint egocentric and exocentric applications.
@@ III Existing Approaches
Some early work has investigated the task of jointly relating egocentric and exocentric vision [48, 49]. In this section, we discuss important tasks jointly modeling vision from the first-person and third-person perspectives.
III-A Identification
@@
It is the task of matching camera wearers in an egocentric video to an exocentric video. The lack of visibility in the egocentric video makes this task challenging. Being able to match a participant in both views is an important preliminary task for joint ego-exo learning. It is a well-researched problem. Ardeshir et al.[50, 51] is one of the first works that proposes a graph-matching technique to solve this problem. Ardeshir et al.[52, 51] further propose a joint approach to tackle temporal alignment and person re-identification. Fig 4 shows some examples of this dataset. Similarly, Han et al. [53] propose a different matching function based on spatial distributions. Fan et al. [54] learns a joint-embedding space. The model proposed by Ardeshir et al. [55] extended to focus on temporal alignment. In Han et al. [56], a conditional random field is proposed to identify the subjects from different viewpoints. Xu et al. [57] perform simultaneous matching and segmentation of the subject across both views.
Figure 4: Ego and corresponding exo-view images taken from Ego2Top [50]
@@
While identification focuses on solely matching the camera wearer, re-identification aims to learn the associations between the different subjects present in the egocentric and exocentric views. Work by Ardeshir et al. [58] is one of the earliest approaches to exploring the task of re-identification between the different views. To enable further research in multi-view video-based re-identification, Basaran et al.[59] release a novel multimodal dataset, consisting of around 176,000176000176,000176 , 000 detections. Han et al. [53] utilizes spatial information such as the view-angle of the camera to perform the association. Han et al. [60] attempts to solve a challenging version of the problem by assuming limited appearance matching and different viewing angles in the ego and exo image. Example images from this dataset can be seen in 5. Han et al.[61] considers another challenging variant of this matching problem having minimal overlap of the field-of-view.
Figure 5: Top and side view-images taken from DMHA dataset released in [60]. Image taken from [60].
III-B Action Recognition
@@
Action Recognition is the task of identifying or assigning a category or multiple categories to the action performed by the subject in the video. The release of GoPro wearable cameras led to a large production of first-person videos. However, limited works have combined ego and exo views for identifying actions. The earliest attempt to recognize human activity across first and multiple third-person cameras was done in [49]. It presents a learnable weighted importance classification approach. Truong et al. [62] learns a geometric constraint to transfer knowledge between the multiple views. Rocha et. al [63] learns an invariant space, based on skeleton pose information. Huang et al. [64] extends to a multi-domain scenario, learning a holographic feature space based on both view-invariant and view-specific features. In Peng et al.[65], virtual features from a first-person perspective are synthesized and combined to perform action recognition. Different from other works, Ramirez et al. [66] incorporates gaze information into the robot’s internal representation for improved imitation of human behavior.
III-C Tracking
@@
Tracking is an important Computer Vision problem, where we estimate the global trajectories and match subjects across the video. Yang et al. [67] is one of the earliest works that jointly identifies and tracks the subjects across the first and third-person views. A deep neural network (DNN), robust to action and motion changes is used to generate the 3D trajectory. Han et al. [68] learns a spatio-temporal correspondence between the images of different viewpoints. 6 shows some sample frames released in their dataset [68]. In a follow-up work, Han et al. [han2021complementary] treats the task as a joint optimization problem. Han et al.[69] extends the optimization approach for relating a single third-person view with multiple first-person view images. Recent work by DivoTrack [70] presents a new baseline for multi-view object tracking. Multi-view tracking has also gained importance in other areas like robotics [71].
Figure 6: Example ego-exo frames taken from the CVMHT dataset released in [68]. Image taken from [68].
III-D Generation
@@
In generation, we aim to synthesize an egocentric image, conditioned on an exocentric image and vice versa. Elfeki et al. [72] was the first landmark dataset for exo-ego synthesis and retrieval. A conditional GAN [73] is used to syntheize first-person images. Refer to Fig. 7 for example frames. Liu et al. [74] also utilizes a variation of a GAN. Similarly, Tang et al. [75, 76] utilize semantic information to generate images in different views. Liu et al. [77] utilize a shared network between the ego-exo frames to aid generation. Liu et al. [78] syntheize egocentric videos by combining the semantic map with GANs. Recent work by Luo et al. [79] presents a diffusion-based technique [80] for exocentric to egocentric video synthesis. Different from all the other works, Luo et al. [81] uses action description and egocentric frames to synthesize a video from the third-person perspective. The new Ego-Exo4D dataset [17] constitutes a benchmark for synthesis.
Figure 7: Pairs from simultaneously recorded Ego-Top and Ego-Side dataset. Image taken from [72].
@@
A lot of progress has been made in a related problem of aerial view to ground view synthesis [82, 83, 84].
III-E Affordance
Figure 8: Examples of frames from Demo2Vec [85].
@@ III-E Affordance
Much attention has been drawn to affordance [86, 87, 88]. The objective is to understand the different possible actions that can be performed with an object. Luo et al. [89, 90] extracts affordance-level features from exocentric human-object interactions and transfers it to the egocentric view. Li et. al [91] extend the same work, but use a weakly-supervised technique. Chen et al. [92] extends affordance learning from videos using an attention-based network. Xu et al. [xu2024weekly] also uses a weakly-supervised technique leveraging cross-view knowledge. Recent work by Zhang et al. [93] integrates a self-explainable module to aid affordance learning. Yang et al. [94] presents a joint coarse and fine-grained feature extraction technique. Different from other techniques, Rai et al. [95] leverage VLM’s knowledge as an auxiliary mask for the task of grounding. Check Fig. 8 for images and corresponding affordance.
III-F Exo-Ego Transfer

### v2

# Egocentric and Exocentric Methods: A Short Survey

@@ Abstract
Egocentric vision captures the scene from the point of view of the camera wearer, while exocentric vision captures the overall scene context. Jointly modeling ego and exo views is crucial to developing next-generation AI agents. The community has regained interest in the field of egocentric vision. While the third-person view and first-person have been thoroughly investigated, very few works aim to study both synchronously. Exocentric videos contain many relevant signals that are transferrable to egocentric videos. This paper provides a timely overview of works combining egocentric and exocentric visions, a very new but promising research topic. We describe in detail the datasets and present a survey of the key applications of ego-exo joint learning, where we identify the most recent advances. With the presentation of the current status of the progress, we believe this short but timely survey will be valuable to the broad video-understanding community, particularly when multi-view modeling is critical.
keywords:
Egocentric , exocentric , ego-exo learning , action recognition.
journal: Computer Vision and Image Understanding
[…]
[inst1]organization=University of Central Florida
[…]
[inst2]organization=Honda Research Institute USA
[…]
[inst3]organization=University of California, Riverside
1 Introduction
Humans perceive the world from multiple viewpoints. We watch Do-it-yourself videos to learn new skills. A bicycle repair video alternates between the ego (first-person) and exo (third-person) viewpoints. An ego (close-up) view of the bicycle captures vital hand-object interactions and an exo (third-person) view captures the overall context in the environment. We can relate the object from the third-person to the first-person perspective. Being able to map skills to one’s own body has been a well-studied problem in cognitive science [1, 2, 3]. Capturing video from both the ego and exo views is a vital frontier for AI to understand human activities. Widespread applications exist in augmented reality [4, 5] and robotics [6, 7, 8].
@@ 1 Introduction
Despite the importance of multi-view learning, most efforts into video understanding have focused on only one view, third-person (exocentric) viewpoint [9, 10, 11, 12, 13, 14] or first-person (egocentric) viewpoints [15] separately. While existing algorithms perform considerably well on third-person settings [16], a significant gap exists in the egocentric settings [17, 18, 19].
An exocentric view contains many relevant cues for recognizing the egocentric view. For example, in Figure 1, the hand-object interaction of “cutting” in the third-person view can be useful to recognize in the first-person view. This short survey provides a timely high-level overview of various egocentric-exocentric learning tasks. Figure 2 provides an overview of these tasks.
The rest of this paper is organized as follows. In Section 2, we introduce the datasets that contain paired ego-exo views. In Section 3, we discuss existing approaches that have recently been presented for various applications. Finally, Section 4 concludes the paper with a summary and provides several promising future research directions.
@@
2 Datasets
Several datasets containing paired ego-exo views have been proposed in the literature [22, 23, 24, 25, 26]. “Mixed” ego-exo views have been covered by [27, 28, 29, 30, 31, 32, 33, 34, 35]. Zhang et al. [36] captures egocentric interactions in a 3D viewpoint. Xue et al.’s AE2 dataset [35] is sampled from existing ego and exo datasets. These datasets have several shortcomings: lack of magnitude, weak synchronization, and poor diversity. The recently released large-scale datasets, such as Ego-Exo4D [17], EgoExoLearn [20], and EgoExo-Fitness [21], attempt to bridge this gap. Please refer to Table 1 for an overview.
Beyond the standard parameters outlined in Table 1, these datasets differ in several key aspects, such as data modalities (e.g., RGB, depth, IMUs - Inertial Measurement Units, stereo), annotation richness (e.g., fine-grained vs. coarse), and the quality of synchronization between the egocentric and exocentic views. While some datasets are specifically designed for action recognition, others prioritize pose estimation or sequence verification. Additionally, datasets vary in their environmental diversity, ranging from controlled lab-settings to real-world outdoor scenarios.
Table 1: Comparison of existing datasets across various parameters, arranged in chronological order.
@@ 2 Datasets
CMU-MMAC dataset [23] is one of the earliest datasets that captures ego and exo video. It is composed of 43 participants cooking 5 recipes in the kitchen setting. Multiple modalities like audio, video, accelerations, and motion capture are present in this dataset.
@@
Charades-Ego dataset [26] was one of the former large-scale joint multi-view dataset efforts containing 68.8 hours of first and third-person video. 112 actors hired by Amazon Mechanical Turk recorded 34 hours of scripted scenarios.
@@
Lemma dataset [37] comprises multi-view and multi-agent daily-life activities. 3D skeletons and RGB-D are collected to give a broad perspective.
@@
Homage dataset [24] is a synchronized multi-view dataset comprising 30 hours of ego-exo video from 27 participants performing household activities in the same environment. It is well annotated with both the hierarchical and atomic action labels.
@@
H2O dataset [22] focuses on 3D egocentric object-level manipulations. It comprises 3D hand poses, 6D object poses, camera poses, object meshes, and scene-point clouds. 4 different participants perform 36 unique actions in three unique environments.
@@
Assembly101 dataset [25] features non-scripted multi-step activities. 101 toy vehicles are manipulated in 4321 video sequences for a total of 513 hours. It constitutes 1380 fine-grained and 202 coarse-grained action classes. AssemblyHands [38] is a subset of Assembly101 to study the challenging problem of 3D hand pose estimation and action classification.
@@
AE2 dataset [35] is one of the premier attempts to learn a view-invariant self-supervised embedding from unpaired ego and exo videos. To this end, they create a new benchmark, sampled from five public datasets [19, 23, 39, 22, 40], and a self-collected tennis dataset. It is composed of 322 clips.
@@
Ego-Exo4D dataset [17] is the largest multi-view dataset including the egocentric view and the corresponding exocentric information. Moreover, it also offers multiple natural language descriptions including expert commentary, narrate-and-act descriptions, and atomic action descriptions. It is rich in modalities like audio, IMUs, video, depth, gaze, stereo, 3D environments, thermal IR, GPS, motion capture, 6DOF, barometer and magnetometer readings. 740 subjects shot 123 scenes across different cities. It releases new challenging benchmarks like keystep recognition, efficient action detection, and proficiency estimation.
@@
EgoExoLearn dataset [20] is another concurrent large-scale ego-exo synchronized dataset. It contains 120 hours of demonstration activities recorded in the lab and daily-life settings. It is richly annotated with fine-grained captions. Unlike previous datasets, it releases benchmarks on cross-view action anticipation and proficiency estimation.
@@
EgoExo-Fitness dataset [21] was also concurrently released along with the previous two ego-exo datasets. While the previous datasets extensively explored daily-life activities, EgoExo-Fitness focuses on exercise-related activities. It comes with a new set of benchmarks for cross-view sequence verification.
3 Existing Approaches
Figure 3: Categorization of the joint egocentric and exocentric tasks. These tasks can be grouped into three major groups: Perception Tasks, Cross-View Learning and Generation. Perception Tasks include identification, tracking, action recognition, and affordance analysis, while Cross-View Learning encompasses ego-exo transfer and joint ego-exo works.
@@ 3 Existing Approaches
Some early work has investigated the task of jointly relating egocentric and exocentric vision [41, 42]. In this section, we discuss important tasks jointly modeling vision from the first-person and third-person perspectives. In the following subsections, we will introduce existing approaches for various joint egocentric and exocentric applications, as illustrated in Figure 3.
3.1 Identification
@@
It is the task of matching camera wearers in an egocentric video to an exocentric video. The lack of visibility in the egocentric video makes this task challenging. Being able to match a participant in both views is an important preliminary task for joint ego-exo learning. It is a well-researched problem. Ardeshir et al. [43, 44] is one of the first works that proposes a graph-matching technique to solve this problem. Ardeshir et al. [44, 45] further propose a joint approach to tackle temporal alignment and person re-identification. Figure 4 presents a unified framework behind these models. Similarly, Han et al. [46] propose a different matching function based on spatial distributions. Fan et al. [47] learns a joint-embedding space. The model proposed by Ardeshir et al. [48] extended to focus on temporal alignment. In Han et al. [49], a conditional random field is proposed to identify the subjects from different viewpoints. Xu et al. [50] perform simultaneous matching and segmentation of the subject across both views.
Figure 4: A modular framework for identification: A feature extraction module (graph/spatial) captures representations from the image inputs and fed into a matching module for identification. Images adapted from [43].
@@
While identification focuses solely on matching the camera wearer, re-identification aims to learn the associations between the different subjects in the egocentric and exocentric views. Work by Ardeshir et al. [51] is one of the earliest approaches to exploring the task of re-identification between the different views. To enable further research in multi-view video-based re-identification, Basaran et al. [52] release a novel multimodal dataset, consisting of around 176,000176000176,000176 , 000 detections. Han et al. [46] utilizes spatial information such as the view-angle of the camera to perform the association. Han et al. [53] attempts to solve a challenging version of the problem by assuming limited appearance matching and different viewing angles in the ego and exo image. Han et al. [54] considers another challenging variant of this matching problem having minimal overlap of the field-of-view.
While current methods assume shared features between ego and exo views, real world cases involve significant distortions and occlusions. Future models should explore occlusion-aware architecture and multi-modal fusion (e.g., depth, IMU sensors) to improve robustness. Further, exploring temporal feature consistency across frames could enhance re-identification accuracy.
Figure 5: A unified framework for joint egocentric-exocentric action recognition. A joint multi-view representation space is learned from the egocentric and exocentric views. At inference, the model utilizes this learned representation to make egocentric video predictions, leveraging knowledge from exocentric views for improved recognition. Images taken from [18] and [55].
3.2 Action Recognition
@@
Action Recognition is the task of identifying or assigning a category or multiple categories to the action performed by the subject in the video. The release of GoPro wearable cameras significantly increased the production of first-person videos. However, limited works have combined ego and exo views for identifying actions. The earliest attempt to recognize human activity across first and multiple third-person cameras was done in [42]. It presents a learnable weighted importance classification approach. Truong et al. [56] learns a geometric constraint to transfer knowledge between the multiple views. Rocha et al. [57] learns an invariant space, based on skeleton pose information. Huang et al. [58] extends to a multi-domain scenario, learning a holographic feature space based on both view-invariant and view-specific features. In Peng et al. [59], virtual features from a first-person perspective are synthesized and combined to perform action recognition. Different from other works, Ramirez et al. [60] incorporates gaze information into the robot’s internal representation for improved imitation of human behavior.
Existing methods rely only on appearance-based features, which are sensitive to occlusions and viewpoint variations. Developing robust multi-modal feature representations (e.g., depth, IMU sensors, etc. can enhance action recognition. Overall, a unified framework of ego-exo action recognition is illustrated in Figure 5.
3.3 Tracking
@@
Tracking is a fundamental Computer Vision problem, that involves estimating the global trajectories and match subjects across the video. Yang et al. [61] is one of the earliest works that jointly identifies and tracks the subjects across the first and third-person views. A deep neural network (DNN), robust to action and motion changes is used to generate the 3D trajectory. Han et al. [62] learns a spatio-temporal correspondence between the images of different viewpoints. In a follow-up work, Han et al. [49] treats the task as a joint optimization problem. Han et al. [63] extends the optimization approach for relating a single third-person view with multiple first-person view images. Recent work by DivoTrack [64] presents a new baseline. CrossMOT for multi-view object tracking, illustrated in Figure 6. Multi-view tracking has also gained importance in other areas like robotics [65].
Most methods rely on manual annotations, making scalability challenging. Future research should explore self-supervised object tracking objectives to mitigate the need for labelled data. Secondly, developing occlusion-aware tracking models is crucial for robustness. Thirdly, current models are computationally expensive and unsuitable for deployment. Finally, multi-modal fusion by incorporating depth sensors and IMUs can enhance the generalization to diverse environments.
Figure 6: Illustration of a tracking framework from [64], showcasing a unified approach for object detection, cross-view identity matching, and single-view tracking.
Figure 7: Illustration of the task that synthesizes egocentric video frames from exocentric video frames, as demonstrated in [66].
3.4 Generation
@@
In generation, we aim to synthesize an egocentric image, conditioned on an exocentric image and vice versa. Elfeki et al. [67] was the first landmark dataset for exo-ego synthesis and retrieval. A conditional GAN [68] is used to synthesize first-person images. Liu et al. [69] also utilizes a variation of a GAN. Similarly, Tang et al. [70, 71] utilize semantic information to generate images in different views. Liu et al. [72] utilize a shared network between the ego-exo frames to aid generation. Liu et al. [73] synthesize egocentric videos by combining the semantic map with GANs. Recent work by Luo et al. [66] presents a diffusion-based approach [74] for synthesizing egocentric videos from exocentric inputs, as illustrated in Figure 7. Different from all the other works, Luo et al. [75] uses action description and egocentric frames to synthesize a video from the third-person perspective. The new Ego-Exo4D dataset [17] constitutes a benchmark for synthesis.
@@
Significant progress has also been made in a related problem of aerial view to ground view synthesis [76, 77, 78].
Despite these advancements, several challenges remain unresolved. Bridging the viewpoint shift between egocentric and exocentric frames continues to be a major obstacle, as current models struggle to generalize across unseen perspectives. Integrating multi-modal signals, such as audio or IMU data, could enhance the realism of generated sequences and improve the scalability and efficiency for real-world deployment. Addressing these challenges will be crucial for advancing the field and enabling more robust, adaptable, and scalable solutions for egocentric-exocentric synthesis.
3.5 Affordance
Figure 8: LOCATE framework [79] utilizes DINO-ViT [80] features for precise localization of human-object interactions.
@@ 3.5 Affordance
Much attention has been drawn to affordance [81, 82, 83]. The objective is to understand the different possible actions that can be performed with an object. Luo et al. [84, 85] extracts affordance-level features from exocentric human-object interactions and transfers it to the egocentric view. Li et al. [79] extend this work using a weakly-supervised technique, as illustrated in Figure 8. Chen et al. [86] extends affordance learning from videos using an attention-based network. Xu et al. [87] also uses a weakly-supervised technique leveraging cross-view knowledge. Recent work by Zhang et al. [88] integrates a self-explainable module to aid affordance learning. Yang et al. [89] presents a joint coarse and fine-grained feature extraction technique. Different from other techniques, Rai et al. [90] leverage knowledge of Vision-Language Models (VLMs) as an auxiliary mask for the task of grounding.
A primary challenge, unaddressed in affordance learning methods is occlusion of critical objects which can obscure important affordance cues. Additionally, variability is introduced due to the diversity of human-object interactions. Finally, integrating multi-modal sensor data, such as depth data and IMUs can provide additional context to affordances. Addressing these challenges will advance the field of affordance learning, particularly in applications involving egocentric and exocentric perspectives.
3.6 Exo-Ego Transfer
Figure 9: Ego-Exo approach for transferring egocentric cues from exocentric videos [91].